# Learning Admissible Heuristics for A*: Theory and Practice

### v1

@@ 3.2 Sample Complexity using Neural Networks
Definition 2 (Neural networks Cheng et al. (, Definition 2.4)).
@@
The bound in Theorem 5 depends on the size of the graph, which may be large. Ideally, we want to leverage neural networks to obtain a tighter bound in terms of the network’s size. The proposed approach from Balcan et al. (2021a) is not guaranteed to remain valid under a neural network’s parameterization. So, we first define a representation function and then derive a bound on $\mathrm{Pdim}(\mathcal{U})$ that accounts for the parameters of the neural network.
@@ 4 Training Framework
PDBs can have imbalanced distributions of states and heuristic values. In the 6‑edge Rubik’s‑Cube PDB, more than $86\%$ of the states fall in classes 7 and 8. Because these classes have large heuristic values, a model that over‑predicts them is likely to violate admissibility on states with lower heuristics. This can be improved using a delta heuristic $h_{\Delta}$ (Sturtevant et al., 2017). Instead of storing distances in a single PDB, we store a small base PDB with a pattern that is a subset of the full PDB and store only the difference between these PDBs $\Delta=h_{\text{large}}-h_{\text{base}}$. At inference time, the final heuristic is reconstructed as $h_{\text{large}}(s)=h_{\text{base}}(s)+\Delta(s)$. Table 3 (Appendix G) shows that subtracting a 4-edge PDB from the 6-edge PDB significantly shifts the class imbalance.
@@ 5 Experiments
In this section, we aim to answer the following key questions: (1) How effective is the proposed training framework in learning strong admissible heuristics? (2) How robust is the proposed loss function to hyperparameter choices? (3) What is the trade-off between model complexity and the strength of the learned heuristic?
@@
To evaluate these questions, we focus on the $3\times 3$ Rubik’s Cube, where, to our knowledge, no previous machine learning methods have learned an admissible heuristic. We selected four PDBs with distinct characteristics: 8-Corner, $\Delta(6,4)$-Edge, 6-Edge, and 7-Edge. All PDBs are sourced from the HOG2 repository [^1] https://github.com/nathansttt/hog2/tree/PDB-refactor. The summary statistics for the PDBs are presented in Table 1, and the complete heuristic distributions are provided in Table 3 (Appendix G). For learning heuristic models, we adopt a neural network architecture based on the ResNet model (He et al., 2016). A complete description of the model architecture and the selected hyperparameters is presented in Appendix G.
Evaluating Learned Heuristics.
The most critical property of a learned heuristic for achieving optimality is its overestimation rate, which should ideally be zero. However, assigning zero to all states, though admissible, offers no useful guidance. Our goal, therefore, is to learn a heuristic that is both admissible and highly informative. Our results are summarized in Table 2. We achieved an overestimation rate which is nearly zero, indicating a fully admissible heuristic, for the 8-corner and $\Delta(6,4)$-Edge PDBs, and only a few thousand overestimated states for the 6-edge and 7-edge PDBs. The CEA loss achieved a fully admissible heuristic in the 8-corner PDB while matching the average heuristic of the original PDB, demonstrating no information loss. Across all PDBs, the overestimation rate for our loss function is nearly $10^{4}\times$ smaller than that obtained with the CE loss. The comparison of the distributions of overestimated states for both loss functions across all PDBs is presented in Figures 58 in Appendix H. These findings suggest that CE loss is ill-suited for learning admissible heuristics in domains satisfying Assumption 4, where the heuristic learning problem can be formulated as a classification task. We attribute the challenges faced by the CE loss in learning admissible heuristics to two main factors: (1) the dataset’s large size and (2) the sparse representation of each state, which includes only the location and rotation of each cubie. In deep learning classification tasks for large datasets (Sun et al., 2017), there are usually no constraints on model size, and the representation for each state is richer.
When compared to the compressed PDB built using the min compression technique, the CEA loss achieves a significantly higher average heuristic in all PDBs, even though a few thousand overestimated states remain in the 6-edge and 7-edge PDBs. This improvement indicates that the CEA loss preserves more information than classical compression techniques such as min compression. To evaluate how far we can reduce model size while maintaining both admissibility and accuracy, we conduct an experiment on the 8-corner PDB (Appendix I). Remarkably, we achieve a $51\times$ compression relative to the original PDB, with performance comparable to the model reported in Table 2.
@@ Appendix A Related Work
There has been significant effort to learn good heuristics for heuristic search algorithms (Chen et al., 2023; Numeroso et al., 2022; Greco et al., 2022; Chen et al., 2024b; Ferber et al., 2020; Yu et al., 2020; Kim & An, 2020; ús Virseda et al., 2013; Heller et al., 2022; Agostinelli et al., 2019; Kirilenko et al., 2023; Chen et al., 2024a; Hao et al., 2024; Bettker et al., 2024). One branch of work uses Graph Neural Networks to better learn the heuristic function by exploiting the graph structure of the problem (Pándy et al., 2022; Shen et al., 2020). In heuristic search problems, the heuristic value of a state can often be approximated from its neighbors’ values, as they differ only by the cost of the connecting edge. Consequently, one can employ bootstrapping methods such as TD-learning (Sutton, 1988) to learn the heuristic value for each state (Ferber et al., 2022; Agostinelli et al., 2024). Veerapaneni et al. (2023) take a different approach than learning the global cost-to-go values, since these can be time-consuming to train and difficult to generalize to new problems. Instead, Veerapaneni et al. (2023) train local heuristics that estimate the cost to escape from small regions for the robot, and later combine these local heuristics with a global heuristic to reduce node expansions. Because data collection for the local heuristic can be slow for large graphs, Veerapaneni et al. (2024) propose an efficient data collection approach that leverages the combinatorial nature of tasks. Another branch of work considers directly learning the policy that decides which node to expand next rather than learning heuristic values (Orseau & Lelis, 2021; Gomoluch et al., 2020; Feng et al., 2022; Choudhury et al., 2018). For instance, Choudhury et al. (2018) train a policy using imitation learning that bases its decisions solely on the portion of the search space uncovered so far.
@@
Garrett et al. (2016) propose focusing on the ordering of states induced by the heuristic, rather than learning exact heuristic values. They employ Rank Support Vector Machines (RankSVM) by using the number of incorrectly ordered pairs of states in each problem as the loss function. Later on, Chrestien et al. (2024) provide theoretical proof that training heuristic functions to produce correct rankings is sufficient for optimal performance, and explain why learning the exact cost-to-go values via mean-squared error regression can be unnecessarily difficult; their experiments on a wide range of tasks also show that ranking-based loss functions outperform regression-based ones. Subsequently, Chrestien et al. (2022) propose the $L^{*}$ loss, which upper-bounds the number of expanded nodes by ensuring that states on the optimal path have lower heuristic values than those off it. Li et al. (2022) and Agostinelli et al. (2021) investigate the admissibility of learned heuristics, guaranteeing full and approximate admissibility, respectively. Núñez-Molina et al. (2024) model the learned heuristic as a Truncated Gaussian, with an admissible heuristic serving as the lower bound. Balcan et al. (2021a) introduce a theoretical framework that provides generalization guarantees for data-driven algorithms by bounding the performance gap between training data and unseen data for algorithms relying on learned parameters. This framework is further applied to establish sample-complexity bounds for general tree-search algorithms (Balcan et al., 2021b; 2022) and for GBFS/A* (Sakaue & Oki, 2022). Cheng et al. address the sample complexity of branch-and-cut problems by modeling the parameter space with neural networks. They also derive a sample complexity bound for learning instance-dependent parameters.
@@ Appendix D Proofs for Heuristic Functions in $\mathbb{R}^{n}$
Between any two of these hyperplanes, the ordering between $v_{i}$ and $v_{j}$—considering all feasible $g(\cdot)$ values—remains unchanged. Above a boundary hyperplane such as $h(v_{i})-h(v_{j})=nc_{0}$, no further hyperplanes are necessary, since $g$ values cannot exceed $nc_{0}$. Consequently, we have in total $2n\cdot\binom{n}{2}$ hyperplanes. By the Shi arrangement (Orlik & Terao, 2013), these hyperplanes partition $\mathbb{R}^{n}$ into $\mathcal{O}\bigl((2n+1)^{n}\bigr)$ regions. Each region corresponds to a unique assignment of heuristic differences for all vertex pairs, which in turn induces a unique ranking of vertices in OPEN, thus a unique performance outcome on any instance $x\in\Pi$. To shatter $N$ instances, we require at least $2^{N}$ such unique $(u_{x_{1}},u_{x_{2}},...,u_{x_{N}})$ tuples; hence, $\mathcal{O}\bigl((2n+1)^{n}\bigr)\;\geq\;2^{N}$. Solving for the largest $N$ in terms of $n$ yields the upper bound $\mathrm{Pdim}(\mathcal{U})=\mathcal{O}(n\log n)$. ∎
@@ Appendix E Expected suboptimality of A* with reopenings
To prove the above theorem, we first state the following lemma from Hart et al. (1968). We omit its proof, which is straightforward and appears in the original work.
@@
Valenzano et al. (2014) showed that for A* without reopening, we have
@@
To apply Proposition 1 to $\hat{\mathcal{U}}$, we require $\Psi_{h}(\cdot):\Pi\to[0,\hat{H}$]. This is not always possible if the heuristic function can assign $h(v)=\infty$ to some vertex $v$. To avoid this, we assume that $\forall v\in V,\;h(v)\neq\infty\;\;\text{if}\;\;h^{*}(v)\neq\infty$, which means the learned heuristic cannot treat any node from which the goal is reachable as a dead-end. We know that $\hat{H}$ is not greater than the upper bound derived on expected suboptimality without reopening. In Valenzano et al. (2014), a worst-case graph and Martelli’s subgraphs (Martelli, 1977) show that without reopenings, the bounds on expected suboptimality are tight. Hence, $\hat{H}$ is not too large relative to suboptimality. For this theorem, measuring the pseudo-dimension of $\mathcal{H}$ directly is challenging, while it is simpler to measure it for the dual class of $\mathcal{H}$. We introduce Assouad’s dual class of $\mathcal{H}$ as follows.
Definition 4 (Assouad (1983)).
@@
$\mathrm{Pdim}(\mathcal{U})=\mathrm{O}\left(\left(\mathrm{Pdim}(\mathcal{F}^{*})+\mathrm{VCdim}(\mathcal{B}^{*})\right)\lg\left(\mathrm{Pdim}(\mathcal{F}^{*})+\mathrm{VCdim}(\mathcal{B}^{*})\right)+\mathrm{VCdim}(\mathcal{B}^{*})\lg K\right).$
@@
Although we introduced a framework with fewer boundary functions, the bound here is not sharper than that of (Sakaue & Oki, 2022). The key limitation is that, even if $\mathcal{F}$ contained only constant functions, the dimension would still be constrained by $\mathrm{VCdim}(\mathcal{B})$. ∎
@@ Appendix F Proof of theorems using neural networks
Lemma 3 (Lemma A.1 in (Cheng et al., )).
@@
Lemma 4 (Theorem A.2 in (Cheng et al., )).
@@
Lemma 5 (Lemma A.3 in (Cheng et al., )).
@@
Lemma 6 (Lemma A.4 in (Cheng et al., )).
@@
The proof of Lemma˜6 is provided in Section 2 of (Bartlett et al., 1998), and the proof of Lemma˜5 is presented in Section A of (Cheng et al., ). ∎

### v2

@@ 3.2 Sample Complexity using Neural Networks
Definition 2 (Neural networks chengsample).
@@
The bound in Theorem 5 depends on the size of the graph, which may be large. Ideally, we want to leverage neural networks to obtain a tighter bound in terms of the network’s size. The proposed approach from Balcan2021-jv is not guaranteed to remain valid under a neural network’s parameterization. So, we first define a representation function and then derive a bound on $\mathrm{Pdim}(\mathcal{U})$ that accounts for the parameters of the neural network.
@@ 4 Training Framework
PDBs can have imbalanced distributions of states and heuristic values. In the 6‑edge Rubik’s‑Cube PDB, more than $86\%$ of the states fall in classes 7 and 8. Because these classes have large heuristic values, a model that over‑predicts them is likely to violate admissibility on states with lower heuristics. This can be improved using a delta heuristic $h_{\Delta}$ (sturtevant2017value). Instead of storing distances in a single PDB, we store a small base PDB with a pattern that is a subset of the full PDB and store only the difference between these PDBs $\Delta=h_{\text{large}}-h_{\text{base}}$. At inference time, the final heuristic is reconstructed as $h_{\text{large}}(s)=h_{\text{base}}(s)+\Delta(s)$. Table 3 (Appendix G) shows that subtracting a 4-edge PDB from the 6-edge PDB significantly shifts the class imbalance.
@@ 5 Experiments
In this section, we aim to answer the following key questions: (1) How effective is the proposed training framework in learning strong admissible heuristics? (2) How robust is the proposed loss function to hyperparameter choices? (3) What is the trade-off between model complexity and the strength of the learned heuristic? (4) How does the bound on generalization error behave as the number of training instances increases?
@@
To evaluate these questions, we focus on the $3\times 3$ Rubik’s Cube, where, to our knowledge, no previous machine learning methods have learned an admissible heuristic. We selected four PDBs with distinct characteristics: 8-Corner, $\Delta(6,4)$-Edge, 6-Edge, and 7-Edge. All PDBs are sourced from the HOG2 repository [^1] https://github.com/nathansttt/hog2/tree/PDB-refactor. The summary statistics for the PDBs are presented in Table 1, and the complete heuristic distributions are provided in Table 3 (Appendix G). For learning heuristic models, we adopt a neural network architecture based on the ResNet model (he2016deep). A complete description of the model architecture and the selected hyperparameters is presented in Appendix G.
Training and Sampling Strategy.
Looking at the PDB heuristic distributions in Table 3 (Appendix G), there is a massive class imbalance in all PDBs. If we uniformly sample training batches, the model overfits the most populated classes. To avoid this, we handle imbalance at the sampling stage. Mini-batches are constructed by uniformly sampling within each heuristic class, and the number of sampled states from each class is proportional to its size. Our setting also differs from standard supervised learning. In a typical setup, we train on a subset and aim to generalize to a small test set that represents a much larger unseen population. Here, each PDB contains all states of a graph, and a ranking function maps every Rubik’s Cube state uniquely to a PDB state. If a model predicts an admissible heuristic for every state in the PDB, we obtain a fully admissible heuristic for Rubik’s Cube. Therefore, the goal is to compress the information in this large dataset into a small model while losing as little information as possible. To this end, we sample training batches from the full dataset for a fixed number of epochs. Many individual states are never seen during training, due to both dataset size and randomness in sampling, but training still reflects the full data distribution.
Post-hoc weight pruning.
One of the constraints in Equation 4 is the model size $\lvert h\rvert$, since fast heuristic evaluation during search is essential. A common approach for improving inference efficiency is post-hoc weight pruning and related compression techniques (han2015deep; micikevicius2017mixed; krishnamoorthi2018quantizing). In our work, we adopt a simple but effective variant: we train all models in 32-bit precision but perform inference in 16-bit precision, substantially reducing latency during search. We also experimented with 8-bit quantized inference; however, the accuracy loss was significant, so we chose not to consider it further.
@@
Evaluating Learned Heuristics.
The most critical property of a learned heuristic for achieving optimality is its overestimation rate, which should ideally be zero. However, assigning zero to all states, though admissible, offers no useful guidance. Our goal, therefore, is to learn a heuristic that is both admissible and highly informative. Our results are summarized in Table 2. We achieved an overestimation rate which is nearly zero, indicating a fully admissible heuristic, for the 8-corner and $\Delta(6,4)$-Edge PDBs, and only a few thousand overestimated states for the 6-edge and 7-edge PDBs. The CEA loss achieved a fully admissible heuristic in the 8-corner PDB while matching the average heuristic of the original PDB, demonstrating no information loss. Across all PDBs, the overestimation rate for our loss function is nearly $10^{4}\times$ smaller than that obtained with the CE loss. The comparison of the distributions of overestimated states for both loss functions across all PDBs is presented in Figures 69 in Appendix H. These findings suggest that CE loss is ill-suited for learning admissible heuristics in domains satisfying Assumption 4, where the heuristic learning problem can be formulated as a classification task. We attribute the challenges faced by the CE loss in learning admissible heuristics to two main factors: (1) the dataset’s large size and (2) the sparse representation of each state, which includes only the location and rotation of each cubie. In deep learning classification tasks for large datasets (sun2017revisiting), there are usually no constraints on model size, and the representation for each state is richer.
Compared to a compressed PDB built with the min–compression technique, CEA achieves a significantly higher average heuristic on all PDBs, even though a few thousand overestimated states remain in the 6-edge and 7-edge cases. This suggests that CEA preserves more information than classical compression methods such as min–compression. For the 8-corner PDB, when the learned model is used in search to assess its strength, the number of nodes generated was less than half of that for the size-matched compressed PDB heuristic (futuhi2025parallel). To evaluate how far we can reduce model size while maintaining admissibility and accuracy, we conduct an experiment on the 8-corner PDB (Appendix I). Remarkably, we achieve a $51\times$ compression relative to the original PDB while maintaining performance comparable to the results in Table 2.
Figure 2: Generalization error vs. training size for the 8-corner PDB.
Empirical analysis of the generalization error.
Up to this point, our experiments have focused on the final overestimation rate of the learned heuristic. Since our sample-complexity analysis establishes a relationship between the generalization error term and the number of training instances, we conduct an additional experiment to empirically examine this behavior. Specifically, we evaluate how the generalization error associated with the inadmissibility functions in Theorem 4 changes with number of training instances. Figure 2 presents the results for 8-corner PDB. We observe that the generalization error decreases as the number of training instances $N$ increases, following a sublinear decay pattern that aligns with the theoretical rate of $\tilde{\mathcal{O}}\!\left(\hat{H}\sqrt{(n+\log(1/\delta))/N}\right)$. This confirms the expected $1/\sqrt{N}$ decay predicted by our sample-complexity analysis.
@@ Appendix A Related Work
There has been significant effort to learn good heuristics for heuristic search algorithms (chen2023goose; numeroso2022learning; greco2022focal; chen2024return; ferber2020neural; yu2020learning; kim2020learning; us2013learning; heller2022neural; agostinelli2019solving; kirilenko2023transpath; chen2024learning; hao2024learned; bettker2024understanding). One branch of work uses Graph Neural Networks to better learn the heuristic function by exploiting the graph structure of the problem (pandy2022learning; shen2020learning). In heuristic search problems, the heuristic value of a state can often be approximated from its neighbors’ values, as they differ only by the cost of the connecting edge. Consequently, one can employ bootstrapping methods such as TD-learning (sutton1988learning) to learn the heuristic value for each state (ferber2022neural; agostinelli2024q). veerapaneni2023learning take a different approach than learning the global cost-to-go values, since these can be time-consuming to train and difficult to generalize to new problems. Instead, veerapaneni2023learning train local heuristics that estimate the cost to escape from small regions for the robot, and later combine these local heuristics with a global heuristic to reduce node expansions. Because data collection for the local heuristic can be slow for large graphs, veerapaneni2024data propose an efficient data collection approach that leverages the combinatorial nature of tasks. Another branch of work considers directly learning the policy that decides which node to expand next rather than learning heuristic values (orseau2021policy; gomoluch2020learning; feng2022left; choudhury2018data). For instance, choudhury2018data train a policy using imitation learning that bases its decisions solely on the portion of the search space uncovered so far.
@@
DBLP:conf/ijcai/GarrettKL16 propose focusing on the ordering of states induced by the heuristic, rather than learning exact heuristic values. They employ Rank Support Vector Machines (RankSVM) by using the number of incorrectly ordered pairs of states in each problem as the loss function. Later on, chrestien2024optimize provide theoretical proof that training heuristic functions to produce correct rankings is sufficient for optimal performance, and explain why learning the exact cost-to-go values via mean-squared error regression can be unnecessarily difficult; their experiments on a wide range of tasks also show that ranking-based loss functions outperform regression-based ones. Subsequently, chrestien2022differentiable propose the $L^{*}$ loss, which upper-bounds the number of expanded nodes by ensuring that states on the optimal path have lower heuristic values than those off it. li2022optimal and agostinelli2021obtaining investigate the admissibility of learned heuristics, guaranteeing full and approximate admissibility, respectively. DBLP:conf/ijcai/Nunez-MolinaAMF24 model the learned heuristic as a Truncated Gaussian, with an admissible heuristic serving as the lower bound. Balcan2021-jv introduce a theoretical framework that provides generalization guarantees for data-driven algorithms by bounding the performance gap between training data and unseen data for algorithms relying on learned parameters. This framework is further applied to establish sample-complexity bounds for general tree-search algorithms (balcan2021sample; balcan2022improved) and for GBFS/A* (sakaue2022sample). chengsample address the sample complexity of branch-and-cut problems by modeling the parameter space with neural networks. They also derive a sample complexity bound for learning instance-dependent parameters.
@@ Appendix D Proofs for Heuristic Functions in $\mathbb{R}^{n}$
Between any two of these hyperplanes, the ordering between $v_{i}$ and $v_{j}$—considering all feasible $g(\cdot)$ values—remains unchanged. Above a boundary hyperplane such as $h(v_{i})-h(v_{j})=nc_{0}$, no further hyperplanes are necessary, since $g$ values cannot exceed $nc_{0}$. Consequently, we have in total $2n\cdot\binom{n}{2}$ hyperplanes. By the Shi arrangement (orlik2013arrangements), these hyperplanes partition $\mathbb{R}^{n}$ into $\mathcal{O}\bigl((2n+1)^{n}\bigr)$ regions. Each region corresponds to a unique assignment of heuristic differences for all vertex pairs, which in turn induces a unique ranking of vertices in OPEN, thus a unique performance outcome on any instance $x\in\Pi$. To shatter $N$ instances, we require at least $2^{N}$ such unique $(u_{x_{1}},u_{x_{2}},...,u_{x_{N}})$ tuples; hence, $\mathcal{O}\bigl((2n+1)^{n}\bigr)\;\geq\;2^{N}$. Solving for the largest $N$ in terms of $n$ yields the upper bound $\mathrm{Pdim}(\mathcal{U})=\mathcal{O}(n\log n)$. ∎
@@ Appendix E Expected suboptimality of A* with reopenings
To prove the above theorem, we first state the following lemma from hart1968formal. We omit its proof, which is straightforward and appears in the original work.
@@
valenzano2014worst showed that for A* without reopening, we have
@@
To apply Proposition 1 to $\hat{\mathcal{U}}$, we require $\Psi_{h}(\cdot):\Pi\to[0,\hat{H}$]. This is not always possible if the heuristic function can assign $h(v)=\infty$ to some vertex $v$. To avoid this, we assume that $\forall v\in V,\;h(v)\neq\infty\;\;\text{if}\;\;h^{*}(v)\neq\infty$, which means the learned heuristic cannot treat any node from which the goal is reachable as a dead-end. We know that $\hat{H}$ is not greater than the upper bound derived on expected suboptimality without reopening. In valenzano2014worst, a worst-case graph and Martelli’s subgraphs (martelli1977complexity) show that without reopenings, the bounds on expected suboptimality are tight. Hence, $\hat{H}$ is not too large relative to suboptimality. For this theorem, measuring the pseudo-dimension of $\mathcal{H}$ directly is challenging, while it is simpler to measure it for the dual class of $\mathcal{H}$. We introduce Assouad’s dual class of $\mathcal{H}$ as follows.
Definition 4 (Assouad1983-vb).
@@
$\mathrm{Pdim}(\mathcal{U})=\mathrm{O}\left(\left(\mathrm{Pdim}(\mathcal{F}^{*})+\mathrm{VCdim}(\mathcal{B}^{*})\right)\log\left(\mathrm{Pdim}(\mathcal{F}^{*})+\mathrm{VCdim}(\mathcal{B}^{*})\right)+\mathrm{VCdim}(\mathcal{B}^{*})\log K\right).$
@@
Although we introduced a framework with fewer boundary functions, the bound here is not sharper than that of (sakaue2022sample). The key limitation is that, even if $\mathcal{F}$ contained only constant functions, the dimension would still be constrained by $\mathrm{VCdim}(\mathcal{B})$. ∎
@@ Appendix F Proof of theorems using neural networks
Lemma 3 (Lemma A.1 in (chengsample)).
@@
Lemma 4 (Theorem A.2 in (chengsample)).
@@
Lemma 5 (Lemma A.3 in (chengsample)).
@@
Lemma 6 (Lemma A.4 in (chengsample)).
@@
The proof of Lemma˜6 is provided in Section 2 of (bartlett1998almost), and the proof of Lemma˜5 is presented in Section A of (chengsample). ∎